# Complex contagions can outperform simple contagions for network reconstruction with dense networks or saturated dynamics

### v1

@@ Abstract
The classical network SIS and threshold models are special cases of this general model [21]. In discrete simple contagions, a node is infected with probability $\beta$ from a single exposure, and exposures are modeled as independent, meaning that the contagion function is $c(\nu,\beta)=1-(1-\beta)^{\nu}$. In complex contagions, multiple exposures are required for the process to spread, and they are most commonly studied using the threshold model [12, 13], where a node adopts an opinion when the number of its neighbors holding that opinion exceeds a critical threshold $\tau$. In that case, the contagion function becomes $c(\nu,\beta,\tau)=\beta\bm{1}{\nu\geq\tau}$, where $\beta=1$. This nonparametric model allows us to define a reconstruction process unrestricted by these two particular cases.
@@
Figure 1: Overview of the problem. (a) A contagion spreads on a network for $T$ time steps, and we observe the resulting sequence of states $\bm{X}$. The probability that a susceptible node (white) becomes infected (red) at the next time step is a function $c(\nu)$ of the number of infected neighbors it has, e.g., $\nu=4$ for the square node highlighted in blue. (b) We compute a nonparametric Bayesian estimate of the contagion function $c(\nu)$. Here, we show an estimate of $c(\nu)$ obtained from a single short realization of the dynamics when the network is known. Error bars show the 50% highest-density posterior interval (HDPI) of $c(\nu)$. (c) We estimate the network and the contagion function $c(\nu)$ simultaneously using the marginals of the posterior distribution, Eq. (7). The reconstruction error goes to 0 as the amount of data $T$ goes to infinity. The shaded regions indicate the 50% HDPI, and lines show the median AUROC across $1,000$ repetitions. (d) The reconstruction quality is determined by the shape of the contagion function, here demonstrated by varying its overall infectivity $\beta$ and the level of complexity $\omega\in[0,1]$. We use the parametrization $c(\nu,\beta,\omega)=(1-\omega)g+\omega h$ where $g(\nu,\beta)=1-(1-\beta)^{\nu}$ describes a simple contagion model, and $h(\nu,\beta)=\beta\bm{1}_{\nu\geq 2}$ describes a complex threshold model.
@@
where $P(\bm{A}\mid\rho)$ is our prior over network structures $\bm{A}$, and where $\rho\in[0,1]$ is a parameter of the network model, a simple Erdös-Rényi model of expected density $\rho$. This model assigns a probability $P(\bm{A}\mid\rho)=\rho^{|E|}(1-\rho)^{\binom{N}{2}-|E|}$ to networks with $|E|$ edges on $N$ nodes. (But see Ref. [9] on replacing this simplistic assumption with more structured models.) Noticing that the parameters of the dynamical process $\theta=(\gamma,\bm{c})$ and the density $\rho$ are all defined on the interval $[0,1]$, we choose independent (conjugate) beta priors with hyperparameters $(a_{\gamma},b_{\gamma})$ for $\gamma$, and likewise for all the other parameters.
@@ Results.
Figure 1 illustrates our nonparametric Bayesian approach using Zachary’s Karate Club network (henceforth ZKC) [22] as an example. We generate synthetic time series of $T=1,000$ states with a recovery rate of $\gamma=0.1$ starting from the initial condition $\bm{x}(0)=\bm{1}$ to prevent premature stochastic extinctions. We study the impact of the contagion function by running these simulations with both the SIS model, defined by $c(\nu,\beta)=1-(1-\beta)^{\nu}$, and the threshold contagion model $c(\nu,\beta,\tau)=\beta\bm{1}{\nu\geq\tau}$ with $\tau=2$ (we also considered $\tau=3$ in the Supporting Information). We find that imperfect reconstruction of the contagion function $\bm{c}$ is possible with short time series, as illustrated in Fig. 1(b). Values of $\nu$ for which we have little data—corresponding to high degrees—are more difficult to estimate accurately, as our nonparametric framework does not benefit from the strong inductive biases of simpler models.
@@
We observe an intermediate value of $T$ where complex contagions reconstruct the network more accurately than simple contagions. In addition, there is a critical quantity of time-series data, after which increasing the amount of data leads to diminishing returns in reconstruction accuracy. Lastly, we construct a mixture of contagion functions parametrized by $\omega\in[0,1]$ (where $\omega=0$ corresponds to a simple contagion and $\omega=1$ to a threshold contagion) Fig. 1(d) demonstrates that network reconstruction depends on both the infectivity of the contagion function and the complexity of the contagion.
Figure 2: The average difference in reconstruction performance between simple and complex contagions for the SIS model and the threshold model with $\tau=2$. Red denotes regions where complex contagions outperform simple contagions. Blue denotes regions where simple contagions outperform complex contagions. The first row visualizes a sample from each network model; the second row compares the performance of the network reconstruction; and the third row compares the estimation of the network density. The gray lines show the basic reproduction number $R_{0}=\beta\sigma/\gamma\in\{1,11,21,...\}$ [25], where $\sigma(\bm{A})$ is the spectral radius of $\bm{A}$, and grow when moving from the lower left corner to the upper right.
Having confirmed that the method works well, we then investigate whether complex or simple contagion models lead to an easier network reconstruction problem. We explore this question through a comprehensive simulation experiment with generative network models that capture structural properties observed in empirical networks. The Erdös-Rényi model [26], which we use to examine the effect of network density, is characterized by $p$, the probability of two nodes connecting at random. We use the network configuration model [27] with a degree distribution $p(k)\propto k^{\alpha}$ supported on $[2,N-1]$ to observe the effects of degree heterogeneity, and we comment that when $\alpha$ is increased, the network density also increases. The clustered network model [28], which we use to represent clique structure, is constructed by forming a bipartite network, where type-1 nodes have degree 2 and type-2 nodes have degree $s$, and then projecting that network onto the type-1 nodes to create cliques of size $s$. The mean degree of the projected network scales with the square of the clique size. We fix the network size to $N=50$ nodes and generated time series of $T=2~{}\times~{}10^{3}$ steps. (We ran the study on two additional models; see the Supporting Information for more details.)
@@
We find structural and dynamical regimes where different types of contagions outperform each other. The regions where simple contagions outperform complex contagions can be sufficiently explained by the basic reproduction number $R_{0}$, calculated as $\beta\sigma/\gamma$, where $\sigma(\bm{A})$ is the spectral radius of $\bm{A}$. For all models, the simple contagion process outperforms the complex contagion process in a region roughly corresponding to $2\leq R_{0}\leq 6$ in the simple contagion. For higher $R_{0}$ values, the complex contagion process outperforms simple contagion as the infection time series saturates. Complex contagion never generates transitions with $\nu=1$ since the threshold forbids it; this allows the algorithm to infer denser substructures without noise introduced by pairwise infection information. Just above this region, there is another region for large enough $R_{0}$ where simple contagion again outperforms complex contagion. This can be understood because complex contagion infers a network close to a complete network, whereas simple contagion infers an empty network—but with high confidence about several of the network links. For the densest of networks, these algorithms perform equally (notice that the densest power law CM is far from a complete network), because there is so much saturation that the elimination of pairwise information is not enough to overcome the confounding noise present when almost every node is infected at once.
There may be cases where, although our algorithm fails to place links correctly, it nonetheless can accurately recover the number of links, i.e., the density, $\rho$. Comparing the performance of complex contagions with respect to simple contagions in the bottom row of Fig. 2, we observe that these regions do not correspond to the performance differences with respect to AUROC. The regions in the upper right of the plots for the Erdös-Rényi and clustered network models where complex contagions vastly outperform simple contagions in estimating the network density are caused by bimodality in the distribution of estimated $\rho$ values inferred from simple contagion time series. This is because our MCMC sampler struggles to sample in this region and gets stuck in local maxima due to the ruggedness of the likelihood landscape. In the lower left of the plots, where the variance in the estimated value of $\rho$ is lower, we see similar behavior as described above for the AUROC.
Figure 3: Difference in reconstruction performance of the network SIS contagion process and the threshold contagion process with threshold $\tau=2$, for nodes of various coreness and dynamical parameters identical to those chosen in Fig. 1(c). This illustrates that for intermediate amounts of infection data, complex contagions outperform simple contagions due to their recovery of $k\geq 2$-core nodes. The shaded regions represent the 50% HDPI, and the lines are the median of 1,000 realizations.
@@ Discussion.
By developing a nonparametric approach to the network inference problem, we avoided model-based biases and were able to study how the complexity of a dynamic process helps or hurts our ability to infer its rules and the network that supports it. We found that complex contagions can have an advantage over equivalent simple contagions if they can avoid saturation and better resolve dense networks.
This result suggests that different dynamical processes have different statistical power to resolve different network structures, with threshold contagions being better in dense cliques when the dynamics are sufficiently intense. More generally, we could imagine optimizing or tuning a dynamic process to infer different network features. Alternatively, it also suggests that different data streams allow us to learn different types of network features depending on the complexity of their generative mechanisms.
Future work should, therefore, explore other network structures and other types of contagions (e.g., relative thresholds, superlinear contagions). In doing so, we should be able to categorize different types of contagions not just by their global accuracy but by how well they can resolve different local features of networks. Our work will not only inspire future network reconstruction methods but also guide how well different types of experimental data can hope to inform network reconstruction.
@@
All code and data used in this study are available on Github [29]. All networks are visualized with XGI [30].
@@
N.W.L., L.H.-D., and J.-G.Y. acknowledge support from the National Institutes of Health 1P20 GM125498-01 Centers of Biomedical Research Excellence Award.

### v3

# Reconstructing networks from simple and complex contagions

@@ Abstract
The classical network SIS and threshold models are special cases of this general model [21]. In discrete simple contagions, a node is infected with probability $\beta$ from a single exposure, and exposures are modeled as independent, meaning that the contagion function is $c(\nu,\beta)=1-(1-\beta)^{\nu}$. In complex contagions, multiple exposures are required for the process to spread, and they are most commonly studied using the threshold model [12, 13], where a node adopts an opinion when the number of its neighbors holding that opinion exceeds a critical threshold $\tau$. In that case, the contagion function becomes $c(\nu,\beta,\tau)=\beta\bm{1}_{\nu\geq\tau}$, where $\beta=1$. This nonparametric model allows us to define a reconstruction process unrestricted by these two particular cases.
@@
Figure 1: Overview of the problem. (a) A contagion spreads on a network for $T$ time steps, and we observe the resulting sequence of states $\bm{X}$. The probability that a susceptible node (white) becomes infected (red) at the next time step is a function $c(\nu)$ of the number of infected neighbors it has, e.g., $\nu=4$ for the square node highlighted in blue. (b) We compute a nonparametric Bayesian estimate of the contagion function $c(\nu)$. Here, we show an estimate of $c(\nu)$ obtained from a single short realization of the dynamics when the network is known. Error bars show the 50% highest-density posterior interval (HDPI) of $c(\nu)$. (c) We estimate the network and the contagion function $c(\nu)$ simultaneously using the marginals of the posterior distribution, Eq. (7). The reconstruction error goes to 0 as the amount of data $T$ goes to infinity. The shaded regions indicate the 50% HDPI, and lines show the median AUROC across $10^{3}$ repetitions. (d) The reconstruction quality is determined by the shape of the contagion function, here demonstrated by varying its overall infectivity $\beta$ and the level of complexity $\omega\in[0,1]$. We use the parametrization $c(\nu,\beta,\omega)=(1-\omega)g+\omega h$ where $g(\nu,\beta)=1-(1-\beta)^{\nu}$ describes a simple contagion model, and $h(\nu,\beta)=\beta\bm{1}_{\nu\geq 2}$ describes a complex threshold model.
@@
where $P(\bm{A}\mid\rho)$ is our prior over network structures $\bm{A}$, and where $\rho\in[0,1]$ is a parameter of the network model, a simple Erdős-Rényi model of expected density $\rho$. This model assigns a probability $P(\bm{A}\mid\rho)=\rho^{|E|}(1-\rho)^{\binom{N}{2}-|E|}$ to networks with $|E|$ edges on $N$ nodes. (But see Ref. [9] on replacing this simplistic assumption with more structured models.) Noticing that the parameters of the dynamical process $\theta=(\gamma,\bm{c})$ and the density $\rho$ are all defined on the interval $[0,1]$, we choose independent (conjugate) beta priors with hyperparameters $(a_{\gamma},b_{\gamma})$ for $\gamma$, and likewise for all the other parameters.
@@ Results.
Figure 1 illustrates our nonparametric Bayesian approach using the Zachary’s Karate Club network (henceforth ZKC) [22] as an example. We generate synthetic time series data comprising nodal states for $T=10^{3}$ time steps with a recovery rate of $\gamma=0.1$ starting from the initial condition $\bm{x}(0)=\bm{1}$ to prevent premature stochastic extinctions. We study the impact of the contagion function by running these simulations with both the SIS model, defined by $c(\nu,\beta)=1-(1-\beta)^{\nu}$, and the threshold contagion model $c(\nu,\beta,\tau)=\beta\bm{1}{\nu\geq\tau}$ with $\tau=2$ (we also considered $\tau=3$ in the Supplemental Material). We find that imperfect reconstruction of the contagion function $\bm{c}$ is possible with short time series, as illustrated in Fig. 1(b). Values of $\nu$ for which we have little data—corresponding to high degrees—are more difficult to estimate accurately, as our nonparametric framework does not benefit from the strong inductive biases of simpler models.
@@
We observe an intermediate value of $T$ where complex contagions reconstruct the network more accurately than simple contagions. In addition, we find that increasing the amount of data leads to diminishing returns in reconstruction accuracy and that the accuracy reaches a plateau close to $1$ for both dynamics in the limit of large $T$. Lastly, we construct a mixture of contagion functions parametrized by $\omega\in[0,1]$ (where $\omega=0$ corresponds to a simple contagion and $\omega=1$ to a threshold contagion) Fig. 1(d) demonstrates that network reconstruction depends on both the infectivity of the contagion function and the complexity of the contagion.
Figure 2: The average difference in reconstruction performance between simple and complex contagions for the SIS model and the threshold model with $\tau=2$. Red denotes regions where complex contagions outperform simple contagions. Blue denotes regions where simple contagions outperform complex contagions. The first row visualizes a sample from each network model; the second row compares the performance of the network reconstruction; and the third row compares the estimation of the network density. The gray lines show the basic reproduction number $R_{0}=\beta\sigma(\bm{A})/\gamma\in\{1,11,21,...\}$ [25], where $\sigma(\bm{A})$ is the spectral radius of $\bm{A}$, and grow when moving from the lower left corner to the upper right.
Having confirmed that the method works well, we then investigate whether complex or simple contagion models lead to an easier network reconstruction problem. We explore this question through a comprehensive simulation experiment with generative network models that capture structural properties observed in empirical networks. The Erdős-Rényi model [26], which we use to examine the effect of network density, is characterized by $p$, the probability of two nodes connecting at random. We use the network configuration model (CM) [27] with a degree distribution $p(k)\propto k^{\alpha}$ supported on $[2,N-1]$ to observe the effects of degree heterogeneity, and we comment that when $\alpha$ is increased, the network density also increases. The clustered network model [28], which we use to represent clique structure, is constructed by forming a bipartite network, where type-1 nodes have degree 2 and type-2 nodes have degree $s$, and then projecting that network onto the type-1 nodes to create cliques of size $s$. The mean degree of the projected network scales with the square of the clique size. We fix the network size to $N=50$ nodes and generated time series of $T=2~{}\times~{}10^{3}$ steps. (We ran the study on two additional models; see the Supplemental Material for more details.)
@@
We find structural and dynamical regimes where different types of contagions outperform each other. The regions where simple contagions outperform complex contagions can be sufficiently explained by the basic reproduction number $R_{0}$, calculated as $\beta\sigma(\bm{A})/\gamma$, where $\sigma(\bm{A})$ is the spectral radius of $\bm{A}$. For all models, the simple contagion process outperforms the complex contagion process in a region roughly corresponding to $2\leq R_{0}\leq 6$ in the simple contagion. For higher $R_{0}$ values, the complex contagion process outperforms simple contagion as the infection time series saturates. Complex contagion never generates transitions with $\nu=1$ since the threshold forbids it; this allows the algorithm to infer denser substructures without noise introduced by pairwise infection information. Just above this region, there is another region for large enough $R_{0}$ where simple contagion again outperforms complex contagion. This can be understood because complex contagion infers a network close to a complete network, whereas simple contagion infers an empty network—but with high confidence about several of the network links. For the densest of networks, these algorithms perform equally (notice that the densest power law CM is far from a complete network), because there is so much saturation that the elimination of pairwise information is not enough to overcome the confounding noise present when almost every node is infected at once.
There may be cases where, although our algorithm fails to place links correctly, it nonetheless can accurately recover the number of links, i.e., the density, $\rho$. Comparing the performance of complex contagions with respect to simple contagions in the bottom row of Fig. 2, we observe that these regions do not correspond to the performance differences with respect to AUROC. The regions in the upper right of the plots for the Erdős-Rényi and clustered network models where complex contagions vastly outperform simple contagions in estimating the network density are caused by bimodality in the distribution of estimated $\rho$ values inferred from simple contagion time series. This is because our MCMC sampler struggles to sample in this region and gets stuck in local maxima due to the ruggedness of the likelihood landscape. In the lower left of the plots, where the variance in the estimated value of $\rho$ is lower, we see similar behavior as described above for the AUROC.
Figure 3: Difference in reconstruction performance of the network SIS contagion process and the threshold contagion process with threshold $\tau=2$, for nodes of various coreness and dynamical parameters identical to those chosen in Fig. 1(c). This illustrates that for intermediate amounts of infection data, complex contagions outperform simple contagions due to their recovery of $k\geq 2$-core nodes. The shaded regions represent the 50% HDPI, and the lines are the median of $10^{3}$ realizations.
@@ Discussion.
By developing a nonparametric approach to the network inference problem, we avoid model-based biases and are able to study how the complexity of a dynamic process helps or hurts our ability to infer its rules and the network that supports it. We find that complex contagions can have an advantage over equivalent simple contagions if they can avoid saturation and better resolve dense networks.
This result suggests that different dynamical processes have different statistical power to resolve different network structures, with threshold contagions being better in dense cliques when the dynamics are sufficiently intense. More generally, we could imagine optimizing or tuning a dynamic process to infer different network features. Future work should, therefore, explore other network structures and other types of contagions (e.g., relative thresholds, superlinear contagions). In doing so, we should be able to categorize different types of contagions not just by their global accuracy but by how well they can resolve different local features of networks.
More practically, it also suggests that different data streams allow us to learn different types of network features depending on the complexity of their generative mechanisms. Our results show that network reconstruction is much more feasible for diseases such as SARS-CoV-2 [29], Mpox [30], or rhinovirus in contrast to highly infectious diseases such as measles [31] or chickenpox [32]. In the context of social media where the spread of information is often modeled as a complex contagion process, our results imply that we will learn less than would be expected if we simply translate prior reconstructability results for simple contagion processes [9] with empirically determined estimates of $R_{0}$ [33]. For extremely viral trends or information, however, we may be able to recover the effective contact networks with more precision than previously estimated. These results will not only inspire future network reconstruction methods but also guide how well different types of experimental data can hope to inform network reconstruction.
@@
All code and data used in this study are available on Github and at Ref. [34]. All networks are visualized with XGI [35].
@@
N.W.L., L.H.-D., and J.-G.Y. acknowledge support from the National Institutes of Health 1P20 GM125498-01 Centers of Biomedical Research Excellence Award. N.W.L. acknowledges support from the University of Virginia Prominence-to-Preeminence (P2PE) STEM Targeted Initiatives Fund, SIF176A Contagion Science.